OpenStudy (anonymous):

Calculate the derivative of the function. g(x) = (2x^2 + x + 1)^−3 Step by step please. I understand you use the chain rule, but then it confuses me after that.

OpenStudy (anonymous):

think of this as \[f(g(x))\] with \[g(x)=2x^2+x+1\] the "inside function" and \[f(u)=u^{-3}\] the "outside function" so that \[f(g(x))=(x^2+x+1)^{-3}\]

OpenStudy (anonymous):

then by the chain rule \[f(g(x))'=f'(g(x))\times g'(x)=-3(2x^2+x+1)^{-4}\times (2x+1)\]

OpenStudy (anonymous):

or think of it like this \[y=u^{-3}\] \[u=2x^2+x+1\] \[\frac{dy}{du}=-3u^{-4}\] \[\frac{du}{dx}=2x+1\] and \[\frac{dy}{dx}=\frac{dy}{du}\times \frac{du}{dx}=-3u^{-4}\times (2x+1)=-3(2x^2+x+1)^{-4}\times (2x+1)\]

OpenStudy (anonymous):

Thank you this was very helpful.

OpenStudy (anonymous):

yw